# Deep learning joint extremes of metocean variables using the SPAR model

### v2

@@ 1 Introduction
Many problems in offshore and coastal engineering require estimation of joint extremes for metocean variables. Responses of offshore and coastal structures are dependent on multiple variables, such as wind speed and direction, wave height, period and direction, flow speed and direction. Providing reliable estimates of the joint extremes in this setting is a challenging problem for metocean engineers. Various design standards recommend the use of the environmental contour method [haver2008environmental]. Some types of contour can be estimated without an explicit model for the joint distribution of variables [Derbanne2019, mackay2023-DIFORM]. However, environmental contour methods typically make simplifying assumptions and only give approximate estimates of long-term extreme responses [haselsteiner2022_NREL, speers2024]. Full probabilistic analysis of long-term extreme responses requires a model for the joint density of the relevant metocean variables. A wide range of approaches have been proposed for estimating joint densities. In the offshore engineering literature, the two most popular approaches are global hierarchical models and copula models – see, e.g., [Ross2020, haselsteiner2021_benchmark].
Let $\mathbf{X}=(X_{1},...,X_{d})\in\mathbb{R}^{d}$, $d\geq 2$ denote a continuous random vector with joint density function $f_{\mathbf{X}}$, and marginal density and cumulative distribution functions $f_{X_{j}}$ and $F_{X_{j}},$ respectively, for $j=1,...,d$. In the global hierarchical approach [bitner2015joint, horn2018new, cheng2019long, simao2022multi], the joint density is written as
@@
where $f_{X_{j}|(X_{1},...,X_{j-1})}$ is the density of $X_{j}$ conditional on $(X_{1},...,X_{j-1})$ for $j\in\{2,\ldots,d\}$. Inference typically involves selecting parametric forms for $f_{X_{1}}$, $f_{X_{2}|X_{1}}$, …, $f_{X_{d}|(X_{1},...,X_{d-1})}$ and estimating relations between the parameters of the conditional densities and the conditioning variables. There are various problems with this approach. Firstly, there is no a priori reason to suppose that variables follow any particular parametric distribution, and misspecified models can have dramatic consequences when approximating dependence structures. Secondly, a model fit to all of the observations does not guarantee a good fit to the tail, which is the region of interest for extremes. Finally, the models for the parameters of the conditional densities are usually based on ad hoc assumptions, and provide no rationale for extrapolating outside the range of observations. In many cases, it has been shown that such models provide a poor fit to observed data, especially in extreme regions [deHauteclocque2022].
@@
where $c:[0,1]^{d}\mapsto[0,\infty)$ is the copula density of $\mathbf{X}$ [Joe2015]. In this case, inference involves choosing parametric models for both the marginal densities and for $c$. As with the global hierarchical approach, there are no a priori reasons to choose particular models. Similarly, fitting to all observations does not guarantee a good fit to the tails. Moreover, different copula models have very different behaviours in the joint tail regions, meaning extrapolation can vary substantially for different choices of copula model [Joe2015].
@@
There are also a wide range of methods in the statistical literature for modelling joint extremes (e.g., [Beirlant2004, Ledford1997, Wadsworth2013]). However, many of these approaches make strong assumptions about the dependence structure, or copula, which are often not supported by environmental datasets [huser2024modeling]. The most popular choice for metocean variables is the conditional extremes model [Heffernan2004], which describes the joint distribution of variables conditional on at least one variable being large. The key limitation of this approach is that it only characterises the region of variable space where the conditioning variable is large, and inferences made using different conditioning variables are not necessarily consistent [Liu2014]. A further limitation of this method (and other methods in the multivariate extremes literature) is that it requires a transformation of the margins to a standard scale. This requires first estimating the marginal distributions for each variable – a process which is subject to uncertainty. Furthermore, it has been demonstrated that poor marginal estimates greatly affect the quality of the resulting multivariate inference [towe2023estimation].
@@
In this paper, we discuss the application of a new method, introduced in [mackay2022imex], which overcomes the limitations of existing approaches and provides a general, flexible framework for modelling multivariate extremes. The model is referred to as the Semi-Parametric Angular-Radial (SPAR) model. The SPAR model provides a framework for estimating multivariate extremes that does not require strong assumptions about the form of the margins or dependence structure, and provides a justified means of extrapolating outside the range of observations. Moreover, the model is only fitted to extreme observations, meaning that no assumptions are required about the bulk of the distribution. Theoretical aspects of the SPAR model are presented in [mackay2023-SPAR], and an inference approach in a two-dimensional setting is provided in [murphy2024inference, mackay2025spar]. The purpose of this paper is to extend the modelling method to the general multivariate setting, with two or more dimensions. The two-dimensional inference scheme in [murphy2024inference, mackay2025spar] utilised cubic regression splines to model the angular dependence of the radial distribution parameters. In this work we take a different approach, and adopt a deep learning scheme, where the radial distribution parameters are modelled using artificial neural networks. As discussed further in Section 3, this offers some computational advantages in higher-dimensional settings.
@@
The paper is organised as follows. Section 2 describes a brief overview of the theoretical aspects of the model. Our deep learning approach for estimating the radial component of the SPAR model is introduced in Section 3. Section 4 presents an example application of the model to a five-dimensional problem: estimating the joint extremes of wind speed, wave direction, wave height, wave period, and wave direction. We discuss the challenges that arise for these particular variables, and how well the model assumptions are satisfied in this setting. Several novel types of diagnostic plots are introduced to assess the fit of the model. We conclude in Section 5 with a discussion and outlook on future work.
@@ 2.1 SPAR model definition
Noting that $\mathbf{X}=R\mathbf{W}$, and that $\mathbf{W}$ lies on the surface of the unit hypersphere, we can see that the ‘extreme’ parts of the distribution of $\mathbf{X}$ correspond to large values of the radial variable at any given angle. Therefore, the problem of modelling multivariate extremes is transformed to that of modelling an angular density $f_{\mathbf{W}}$ and the tail of the conditional radial density $f_{R|\mathbf{W}}$. For a given angle $\mathbf{w}$, the density $f_{R|\mathbf{W}}(r|\mathbf{w})$ is univariate. Univariate extreme value theory suggests that a suitable model for the tail of $f_{R|\mathbf{W}}$ is the GP distribution, with parameters conditional on angle (e.g., [Coles2001]). This motivates the SPAR model, whereby parametric and non-parametric models are used to model the conditional radial and angular distributions, respectively. Define a threshold function $u(\mathbf{w})>0$ to be the quantile of $R|(\mathbf{W}=\mathbf{w})$ with exceedance probability $\zeta\in(0,1)$, with $\zeta$ close to 0, i.e., the solution of $\zeta=\Pr(R>u(\mathbf{w})|\mathbf{W}=\mathbf{w})$. The SPAR model can be written as
@@ 2.2 Exceedance probability contours
As described in [mackay2025spar], the SPAR model provides an explicit means for calculating a contour with a specified exceedance probability $\beta\in(0,\zeta]$. The radius of the contour at angle $\mathbf{w}$ is simply the quantile of the GP distribution at exceedance probability $\beta/\zeta$, given by
@@
This contour is defined in terms of the probability of an observation falling anywhere outside the contour region, or the ‘total exceedance probability’. As such, these contours are more conservative than those defined in terms of marginal exceedance probabilities, such as IFORM contours (or variants thereof), with the conservatism increasing with the number of dimensions [mackay2021marginal]. Moreover, if the primary interest of the analysis is to estimate environmental contours, then the use of the SPAR model is not necessary. Instead, we recommend the use of the Direct-IFORM method [Derbanne2019, mackay2023-DIFORM], which does not require a model for the joint density or any assumptions about the dependence structure between the variables.
The contours defined in (9) can be projected into two dimensions. In Cartesian space, points on the contour are given by $\mathbf{x}=r_{\beta}(\mathbf{w})\,\mathbf{w}$ for $\mathbf{w}\in\mathbb{S}^{d-1}$. The radii $r_{\beta}\geq 0$ can be computed for a discrete set of points on the sphere (see Appendix A), and the projection into dimensions $i,j\in\{1,...,d\}$ is just the $i$ and $j$ components of $\mathbf{x}$.
However, in our application it was found that these contours do not provide a useful diagnostic tool. This is because they do not account for the angular density. In regions of low angular density, estimates of the contours can be highly uncertain. Simpson and Tawn [simpson2024inference] proposed to mitigate for this by either setting the radius to zero when the angular density is zero, or adjusting the exceedance level to account for the angular density. However, the latter approach requires an integral over angles that becomes impractical in higher dimensions. We have therefore opted to use alternative diagnostics, described in Section 4.5.
@@ 3 Inference
Inference for the SPAR model involves estimating the angular density, and the GP threshold and parameter functions. Since the angular-radial density factorises (see (5)), these problems are separable: inference for $f_{\mathbf{W}}$ can be conducted independently of that for $(u(\mathbf{w}),\sigma(\mathbf{w}),\xi(\mathbf{w}))$. Inference for the angular density is discussed in Section 3.1, and modelling of the conditional radial variable is discussed in Section 3.2. Code for fitting our model is available upon reasonable request.
@@ 3.1 Angular modelling
Estimation of densities on the hypersphere $\mathbb{S}^{d-1}$ is part of a discipline known as directional statistics [mardia2000, ley2017modern]. The key difference from estimation of densities on $\mathbb{R}^{d}$ is that the surface of the hypersphere is periodic and bounded, and so distributions defined on $\mathbb{S}^{d-1}$ must conserve these constraints. Various parametric and non-parametric approaches have been developed for estimating densities on the hypersphere, which are directly analogous to approaches used in Euclidean space. These include kernel density (KD) estimation [hall1987kernel, bai1989kernel], mixture models [mooney2003fitting, fu2008modified, hornik2014movmf], and spline-based methods [ferreira2008directional]. See [Pewsey2021] for a recent review of non-parametric approaches, and [wessel2025comparison] for a comparison of various generative deep learning methods for modelling angular densities.
@@ 3.1.2 Choice of kernel
The choice of kernel function has a relatively minor impact on the accuracy of the density estimate compared to the choice of bandwidth. While different kernels can slightly alter the shape of the density estimate, the bandwidth parameter, which controls the amount of smoothing, is the most crucial factor affecting the accuracy [scott2015multivariate]. Here we are mainly guided by computational considerations. The most common choice of kernel for KD models on the sphere is the von Mises-Fisher (vMF) distribution [fisher1953dispersion]. However, simulation from this distribution requires rejection sampling, which can be slow. Therefore, we opt to use the power spherical (PS) distribution [deCao2020psd]. The PS and vMF distributions are both rotationally symmetric distribution about the mean direction, and have similar shapes. However, simulation from the PS distribution is much faster, as discussed further in Appendix B. For $\mathbf{w}\in\mathbb{S}^{d-1}$ the density of the PS distribution is given by
@@ 3.1.3 Bandwidth optimisation
There are various ways in which the bandwidth parameter for the KD model can be optimised, as discussed in [ley2017modern]. Here we use a cross-validation scheme to minimise the negative log-likelihood (NLL) for a hold-out sample. This is the most computationally expensive part of the method – for each test point, the kernel must be evaluated for each observation in the training sample. For a leave-one-out cross-validation scheme, this entails $(n-1)^{2}$ evaluations of the kernel:
@@
where $\hat{f}_{\neq i}(\mathbf{w}_{i})$ is the KD estimate (10) of the density from all observations except $\mathbf{w}_{i}$, evaluated at $\mathbf{w}_{i}$ (i.e., the predictive likelihood). However, as the NLL is effectively an expectation over $\mathbf{W}$, we can take a stochastic approach and estimate the NLL from a smaller sample of size $m$, chosen at random from the observations, so that we only need $(n-1)m$ kernel evaluations. This is similar to the approach used in stochastic gradient descent (see below) in machine learning, where a small batch of observations is used to approximate the true gradient of the cost function.
@@
[…]
@@ 3.1.4 Simulation from a kernel density estimate
Once the value of the bandwidth parameter has been optimised, it is straightforward to simulate from the KD model. As mentioned above, the KD model is effectively a mixture model, with one component corresponding to each observation, centred at the location of the observation. Therefore, to generate a random sample from the KD model requires two steps. First, we sample an observed angle $\mathbf{w}_{i}$ at random (i.e., a random integer $i\in\{1,...,n\}$). This corresponds to choosing one component from the mixture model at random. Secondly, we generate a random value from the kernel with mean direction $\mathbf{w}_{i}$ and bandwidth $\kappa$. As described in Appendix B, simulation from the PS kernel on $\mathbb{S}^{d-1}$ requires generating a beta random variable and $d-1$ standard normal random variables, and applying a rotation. Computationally this is very efficient. For the example presented in Section 4, generating a sample of size $2.7\times 10^{7}$ (100 times the observed sample size) took approximately 10 s on a laptop with specifications mentioned above.

### v3

@@ 1 Introduction
Many problems in offshore and coastal engineering require estimation of joint extremes for metocean variables. Responses of offshore and coastal structures are dependent on multiple variables, such as wind speed and direction, wave height, period and direction, flow speed and direction. Providing reliable estimates of the joint extremes in this setting is a challenging problem for metocean engineers. Various design standards recommend the use of the environmental contour method [1]. Some types of contour can be estimated without an explicit model for the joint distribution of variables [2, 3]. However, environmental contour methods typically make simplifying assumptions and only give approximate estimates of long-term extreme responses [4, 5]. Full probabilistic analysis of long-term extreme responses requires a model for the joint density of the relevant metocean variables. A wide range of approaches have been proposed for estimating joint densities. In the offshore engineering literature, the two most popular approaches are global hierarchical models and copula models – see, e.g., [6, 7].
Let $\mathbf{X}=(X_{1},...,X_{d})\in\mathbb{R}^{d}$, $d\geq 2$ denote a continuous random vector with joint density function $f_{\mathbf{X}}$, and marginal density and cumulative distribution functions $f_{X_{j}}$ and $F_{X_{j}},$ respectively, for j=1,,dj=1,...,d. In the global hierarchical approach [8, 9, 10, 11], the joint density is written as
@@
where $f_{X_{j}|(X_{1},...,X_{j-1})}$ is the density of $X_{j}$ conditional on $(X_{1},...,X_{j-1})$ for $j\in\{2,\ldots,d\}$. Inference typically involves selecting parametric forms for $f_{X_{1}}$, $f_{X_{2}|X_{1}}$, …, $f_{X_{d}|(X_{1},...,X_{d-1})}$ and estimating relations between the parameters of the conditional densities and the conditioning variables. There are various problems with this approach. Firstly, there is no a priori reason to suppose that variables follow any particular parametric distribution, and misspecified models can have dramatic consequences when approximating dependence structures. Secondly, a model fit to all of the observations does not guarantee a good fit to the tail, which is the region of interest for extremes. Finally, the models for the parameters of the conditional densities are usually based on ad hoc assumptions, and provide no rationale for extrapolating outside the range of observations. In many cases, it has been shown that such models provide a poor fit to observed data, especially in extreme regions [12].
@@
where $c:[0,1]^{d}\mapsto[0,\infty)$ is the copula density of $\mathbf{X}$ [13]. In this case, inference involves choosing parametric models for both the marginal densities and for $c$. As with the global hierarchical approach, there are no a priori reasons to choose particular models. Similarly, fitting to all observations does not guarantee a good fit to the tails. Moreover, different copula models have very different behaviours in the joint tail regions, meaning extrapolation can vary substantially for different choices of copula model [13].
@@
There are also a wide range of methods in the statistical literature for modelling joint extremes (e.g., [14, 15, 16]). However, many of these approaches make strong assumptions about the dependence structure, or copula, which are often not supported by environmental datasets [17]. The most popular choice for metocean variables is the conditional extremes model [18], which describes the joint distribution of variables conditional on at least one variable being large. The key limitation of this approach is that it only characterises the region of variable space where the conditioning variable is large, and inferences made using different conditioning variables are not necessarily consistent [19]. A further limitation of this method (and other methods in the multivariate extremes literature) is that it requires a transformation of the margins to a standard scale. This requires first estimating the marginal distributions for each variable – a process which is subject to uncertainty. Furthermore, it has been demonstrated that poor marginal estimates greatly affect the quality of the resulting multivariate inference [20].
@@
In this paper, we discuss the application of a new method, introduced in [21], which overcomes the limitations of existing approaches and provides a general, flexible framework for modelling multivariate extremes. The model is referred to as the Semi-Parametric Angular-Radial (SPAR) model. The SPAR model provides a framework for estimating multivariate extremes that does not require strong assumptions about the form of the margins or dependence structure, and provides a justified means of extrapolating outside the range of observations. Moreover, the model is only fitted to extreme observations, meaning that no assumptions are required about the bulk of the distribution. Theoretical aspects of the SPAR model are presented in [22], and an inference approach in a two-dimensional setting is provided in [23, 24]. The purpose of this paper is to extend the modelling method to the general multivariate setting, with two or more dimensions. The two-dimensional inference scheme in [23, 24] utilised cubic regression splines to model the angular dependence of the radial distribution parameters. In this work we take a different approach, and adopt a deep learning scheme, where the radial distribution parameters are modelled using artificial neural networks. As discussed further in Section 3, this offers some computational advantages in higher-dimensional settings.
@@
The paper is organised as follows. Section 2 describes a brief overview of the theoretical aspects of the model. Our deep learning approach for estimating the radial component of the SPAR model is introduced in Section 3. Section 4 presents an example application of the model to a five-dimensional problem: estimating the joint extremes of wind speed, wave direction, wave height, wave period, and wave direction. We discuss the challenges that arise for these particular variables, and how well the model assumptions are satisfied in this setting. Several novel types of diagnostic plots are introduced to assess the fit of the model. A discussion and outlook on future work is presented in Section 5. Finally, conclusions are presented Section 6 .
@@ 2.1 SPAR model definition
Noting that $\mathbf{X}=R\mathbf{W}$, and that $\mathbf{W}$ lies on the surface of the unit hypersphere, we can see that the ‘extreme’ parts of the distribution of $\mathbf{X}$ correspond to large values of the radial variable at any given angle. Therefore, the problem of modelling multivariate extremes is transformed to that of modelling an angular density $f_{\mathbf{W}}$ and the tail of the conditional radial density $f_{R|\mathbf{W}}$. For a given angle $\mathbf{w}$, the density $f_{R|\mathbf{W}}(r|\mathbf{w})$ is univariate. Univariate extreme value theory suggests that a suitable model for the tail of $f_{R|\mathbf{W}}$ is the GP distribution, with parameters conditional on angle (e.g., [25]). This motivates the SPAR model, whereby parametric and non-parametric models are used to model the conditional radial and angular distributions, respectively. Define a threshold function $u(\mathbf{w})>0$ to be the quantile of $R|(\mathbf{W}=\mathbf{w})$ with exceedance probability $\zeta\in(0,1)$, with $\zeta$ close to 0, i.e., the solution of $\zeta=\Pr(R>u(\mathbf{w})|\mathbf{W}=\mathbf{w})$. The SPAR model can be written as
@@ 2.2 Exceedance probability contours
As described in [24], the SPAR model provides an explicit means for calculating a contour with a specified exceedance probability $\beta\in(0,\zeta]$. The radius of the contour at angle $\mathbf{w}$ is simply the quantile of the GP distribution at exceedance probability $\beta/\zeta$, given by
@@
This contour is defined in terms of the probability of an observation falling anywhere outside the contour region, or the ‘total exceedance probability’. As such, these contours are more conservative than those defined in terms of marginal exceedance probabilities, such as IFORM contours (or variants thereof), with the conservatism increasing with the number of dimensions [26]. Moreover, if the primary interest of the analysis is to estimate environmental contours, then the use of the SPAR model is not necessary. Instead, we recommend the use of the Direct-IFORM method [2, 3], which does not require a model for the joint density or any assumptions about the dependence structure between the variables.
The contours defined in (9) can be projected into two dimensions. In Cartesian space, points on the contour are given by $\mathbf{x}=r_{\beta}(\mathbf{w})\,\mathbf{w}$ for $\mathbf{w}\in\mathbb{S}^{d-1}$. The radii $r_{\beta}\geq 0$ can be computed for a discrete set of points on the sphere (see Appendix A), and the projection into dimensions $i,j\in\{1,...,d\}$ is just the $i$ and $j$ components of $\mathbf{x}$. However, in our application it was found that these contours do not provide a useful diagnostic tool. This is because they do not account for the angular density. In regions of low angular density, estimates of the contours can be highly uncertain. Simpson and Tawn [27] proposed to mitigate for this by either setting the radius to zero when the angular density is zero, or adjusting the exceedance level to account for the angular density. However, the latter approach requires an integral over angles that becomes impractical in higher dimensions. We have therefore opted to use alternative diagnostics, described in Section 4.5.
@@ 3 Inference
Inference for the SPAR model involves estimating the angular density, and the GP threshold and parameter functions. Since the angular-radial density factorises (see (5)), these problems are separable: inference for $f_{\mathbf{W}}$ can be conducted independently of that for $(u(\mathbf{w}),\sigma(\mathbf{w}),\xi(\mathbf{w}))$. Inference for the angular density is discussed in Section 3.1, and modelling of the conditional radial variable is discussed in Section 3.2. Code for fitting our model is available at https://github.com/callumbarltrop/DeepSPAR.
@@ 3.1 Angular modelling
In the SPAR method we estimate the angular density of all observations – see (6). This is different to ‘classical’ multivariate extreme value modelling, where it is only the angular density of ‘extreme’ observations that is modelled (see e.g. [28]. Using the entire sample gives more observations to work with, reducing sampling uncertainties. However, in contrast classical multivariate extremes, in the SPAR approach there are no theoretically-motivated parametric models for the angular density, meaning that flexible modelling approaches are required.
Estimation of densities on the hypersphere $\mathbb{S}^{d-1}$ is part of a discipline known as directional statistics [29, 30]. The key difference from estimation of densities on $\mathbb{R}^{d}$ is that the surface of the hypersphere is periodic and bounded, and so distributions defined on $\mathbb{S}^{d-1}$ must conserve these constraints. Various parametric and non-parametric approaches have been developed for estimating densities on the hypersphere, which are directly analogous to approaches used in Euclidean space. These include kernel density (KD) estimation [31, 32], mixture models [33, 34, 35], and spline-based methods [36]. See [37] for a recent review of non-parametric approaches, and [38] for a comparison of various generative deep learning methods for modelling angular densities.
@@ 3.1.2 Choice of kernel
The choice of kernel function (amongst commonly-used parametric forms for angular densities) has a relatively minor impact on the accuracy of the density estimate compared to the choice of bandwidth. While different kernels can slightly alter the shape of the density estimate, the bandwidth parameter, which controls the amount of smoothing, is the most crucial factor affecting the accuracy [39]. Here we are mainly guided by computational considerations. The most common choice of kernel for KD models on the sphere is the von Mises-Fisher (vMF) distribution [40]. However, simulation from this distribution requires rejection sampling, which can be slow. Therefore, we opt to use the power spherical (PS) distribution [41]. The PS and vMF distributions are both rotationally symmetric distribution about the mean direction, and have similar shapes. However, simulation from the PS distribution is much faster, as discussed further in Appendix B. For $\mathbf{w}\in\mathbb{S}^{d-1}$ the density of the PS distribution is given by
@@ 3.1.3 Bandwidth optimisation
There are various ways in which the bandwidth parameter for the KD model can be optimised, as discussed in [30]. Here we use a cross-validation scheme to minimise the negative log-likelihood (NLL) for a hold-out sample. This is the most computationally expensive part of the method – for each test point, the kernel must be evaluated for each observation in the training sample. For a leave-one-out cross-validation scheme, this entails $(n-1)^{2}$ evaluations of the kernel:
@@
where $\hat{f}_{\neq i}(\mathbf{w}_{i})$ is the KD estimate (10) of the density from all observations except $\mathbf{w}_{i}$, evaluated at $\mathbf{w}_{i}$ (i.e., the predictive likelihood). However, as the NLL is effectively an expectation over $\mathbf{W}$, we can take a stochastic approach and estimate the NLL from a smaller sample of size $m$, chosen at random from the observations, so that we only need (n1)m(n-1)m kernel evaluations. This is similar to the approach used in stochastic gradient descent (see below) in machine learning, where a small batch of observations is used to approximate the true gradient of the cost function.
@@
In the example presented in Section 4, we found that using $m=1000$ prediction points for estimating the NLL gave smooth and repeatable results for our sample size of $n=271,704$ hourly observations, requiring $\approx 271\times 10^{6}$ evaluations of the kernel (a speed up of $\approx 271$ times compared to the full cross-validation scheme). We also set $k=48$, corresponding to a two-day exclusion around each point used to calculate the predictive NLL. Computations of the NLL for 50 logarithmically-spaced test values of $\kappa\in[10^{1},\,10^{4}]$ took approximately 5 minutes on a laptop with an Intel Core i7-1355U 1.7GHz processor. The results are shown in Figure 1, which indicates an optimal value of $\kappa\approx 1200$.
@@ 3.1.4 Simulation from a kernel density estimate
Once the value of the bandwidth parameter has been optimised, it is straightforward to simulate from the KD model. As mentioned above, the KD model is effectively a mixture model, with one component corresponding to each observation, centred at the location of the observation. Therefore, to generate a random sample from the KD model requires two steps. First, we sample an observed angle $\mathbf{w}_{i}$ at random (i.e., a random integer $i\in\{1,...,n\}$). This corresponds to choosing one component from the mixture model at random. Secondly, we generate a random value from the kernel with mean direction $\mathbf{w}_{i}$ and bandwidth $\kappa$. As described in Appendix B, simulation from the PS kernel on $\mathbb{S}^{d-1}$ requires generating a beta random variable and d1d-1 standard normal random variables, and applying a rotation. Computationally this is very efficient. For the example presented in Section 4, generating a sample of size $2.7\times 10^{7}$ (100 times the observed sample size) took approximately 10 s on a laptop with specifications mentioned above.